OpenStudy (anonymous):

Find a formula for Sum of the first n terms of the series- a) √3 + 3 + 3√3 +9.... b) 0.9 + 0.09 + 0.009....

hartnn (hartnn):

familiar with the formula of sum of infinite terms in an Geometric sequence??

OpenStudy (shubhamsrg):

i dont think it goes to infinty!

OpenStudy (shubhamsrg):

infinity*

hartnn (hartnn):

oops, n terms.....so ther's a formula for sum of 'n' terms of GP,familiar with that ? @amishra

OpenStudy (anonymous):

Yes, I got the formula, but I didn't get the answer

hartnn (hartnn):

whats your r in both case.....common ratio??

OpenStudy (anonymous):

For the first one, it's √3, and for the second, its 0.1.

hartnn (hartnn):

correct...so after substituting in that formula , u get \[\frac{\sqrt{3}(-1+(\sqrt{3})^n)}{(-1+\sqrt{3})}\] and \[\frac{0.1(1-(0.1)^n)}{0.9}\] are these the answers u get or the given answers?

OpenStudy (anonymous):

Yes, I got these answers, but the given answers are more simplified :/

hartnn (hartnn):

like ? can u show them....then i'll try to simplify

OpenStudy (anonymous):

a) (3+√3) * ((√3)^n -1) / 2 b) 1- (0.1)^n

hartnn (hartnn):

in the 2nd equation,the numerator should have 0.9 instead of 0.1(as 1st term=0.9) so the 2nd equation directly simplifies to 1-(0.1)^n.....because 0.9 cancels in numerator and denominator

OpenStudy (anonymous):

Oh, yeah, I got that, what about the first one?

hartnn (hartnn):

for 1st they have rationalized the denominator by multiplying (1+root 3) in numerator and denominator.....so denominator became (3-1)=2...got it?

OpenStudy (anonymous):

Oh! I got it now! Wow! Thank you so much, again!!! :D I would have died if you weren't there! :D Thank you!